I need help for this problem. First up, I cant seem to remember what the problem wants me to look for. And I also need to know the answer for that missing question. Here it is: "If the area of the triangle is 240 square meters and one leg is 8 more than the other leg, find the (*)?" THANKS.
Answer by Theo(13342)

1/2 b*h = 240
b = h+8

formula becomes:
1/2 * (h+8)*h = 240
multiply both sides of this equation by 2 to get:
(h+8) * h = 480
remove parentheses to get:
h^2 + 8h = 480
subtract 480 from both sides of the equation to get:
h^2 + 8h - 480 = 0
Solve this equation using the quadratic formula to get:
h = 18.27105745
b = h+8 = 26.27105745

1/2*b*h = 240 which proves the values for b and h are good.

I believe they are asking you for the other leg, which is b.

I don't know what else they would be asking for.